carmelle:

What is the greatest common factor of x^4 and x^3? x x^2 x^3 x^4

AZ:

Let's look at the factors of x^4 and x^3 x^4 = x* x * x * x x^3 = x * x * x how many `x`'s are common to x^4 and x^3

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Let's look at the factors of x^4 and x^3 x^4 = x* x * x * x x^3 = x * x * x how many `x`'s are common to x^4 and x^3 \(\color{#0cbb34}{\text{End of Quote}}\) 3 I think

AZ:

That's right!! now what's another way to write x * x * x

carmelle:

x^3

AZ:

And that right there is your greatest common factor!

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken got to be x^3 \(\color{#0cbb34}{\text{End of Quote}}\) no direct answers

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ And that right there is your greatest common factor! \(\color{#0cbb34}{\text{End of Quote}}\) tysm

AZ:

you're welcome!